We need 15 quarts of 30% antifreeze solution. We have 20% antifreeze and 50% antifreeze to mix together to make it. How much of each should we mix together?
Answer by nerdybill(7384)

We need 15 quarts of 30% antifreeze solution. We have 20% antifreeze and 50% antifreeze to mix together to make it. How much of each should we mix together?
.
Let x = amount (quarts) of 20%
then
15-x = amount (quarts) of 50%
.
.20x + .50(15-x) = .30(15)
.20x + 7.5-.50x = 4.5
7.5-.30x = 4.5
-.30x = -3
x = 10 quarts (20%)
.
Amount of 50%:
15-x = 15-10 = 5 quarts